Intro Long Gone Corp. had total operating expenses of $84 million last year, including depreciation, and paid $5.2 million inIB Attempt 1/10 for 10 pts. Part 3 What was the equity multiplier? k+ decimals Submit
Answer #1

Dividend Payout Ratio = Dividends Paid/Earnings available to shareholders

0.2 = 54000000/Earnings available to shareholders

Therefore, Earnings available to equity shareholders = 54000000/0.2 = 270000000

Retained Earnings = Earnings available to shareholders - Dividends Paid = 270000000-54000000 = 216000000

Addition to retained earnings increased equity by 50%

Therefore, Opening Equity = 2*Retained Earnings = 2*216000000 = 432000000

Debt = Interest Paid/Interest Rate on Debt = 5200000/0.065 = 80000000

From above,

Total Equity = Opening Equity + Current Retained Earnings = 432000000+216000000 = 648000000

Total Assets(assuming no current assests or liabilities) = Total Equity + Debt = 648000000+80000000 = 728000000

Equity Multiplier = Total Assets/Total Equity = 728000000/648000000 = 1.12

(If this was helpful then please rate positively. Thank You:)